I. Show necessary solutions/explanations to support your answer. 

1. How much is the interest if you loan ₱ 350,000 for one year at a simple interest rate of 8%?

2. Which of the following is also referred as Banker’s rule? Explain your answer.

a. exact interest exact time

b. exact interest approximate time

c. ordinary interest exact time

d. ordinary interest approximate time

e. none of these

3. What is the approximate time from April 8, 1999 to October 25, 1999?

4. How many ₱500 monthly payments can be withdrawn from an account containing 

₱35,000 and invested at 8% converted monthly. The first withdraw takes place in one 

Month.

5. On November 15, 2020, Jose Boliva went to BDO Bank and made a loan of ₱100,000

At 10% ordinary interest for 90 days. What is the maturity value of the loan?



1
Expert's answer
2021-04-29T17:19:14-0400

You


1)"I=P*R*T"

I=interest

P=principal

R=rate

T=time

"=350000\times \frac{8}{100}\times1"

"=28000"


2)c. ordinary interest exact time.

Bankers rule considers one year to have 360 days and any fractional Part of the year is the exact number of the loan.


3) Approximate time assumes that each month to have 30 days.

April="(30-8)=22"

May, June, July, August, September="(5\times30)=150"

October="25"

="22+150+25=197days"

 A week has 7days

"\frac{197}{7}=" 28 weeks and 1day

1day has 24 hours

"197\times24=4728" hours


4) "\frac{35000}{500}=70"


"\frac{70}{0.08}=875"


5) "MV=P\times(1+r)^{n}"

MV=maturity value

P=principal

r=interest rate

n=number of compounding intervals

"=100000\times(1+\frac{10}{100})^{\frac{90}{360}}"


"=100000\times(1.1)^{0.25}"

"=100000\times1.024113689"

"=102,411.3689"